OpenStudy (anonymous):

The area of a rectangle is 60 in.2 and the width is 4 in. What is the rectangle’s length @Opcode my answer :l=15 in other choices: 26 inches, or 56 inches

OpenStudy (opcode):

Area of a rectangle: \[A = l \times w\] \[60 = l \times 4\] \[60 \div 4 =l\] \[15 = l\]

OpenStudy (anonymous):

yayy ;D wAit so I am right?

OpenStudy (opcode):

You said 15 right? If you did then you are correct. :-)

OpenStudy (anonymous):

yayy, theres afew left but I think I might have them right but I wanna make sure if that's okay?

OpenStudy (opcode):

Sure.

OpenStudy (anonymous):

The perimeter of a rectangle is 36 yd and the length is 15 yd. What is the rectangle’s width 3yd?

OpenStudy (opcode):

Perimeter of a rectangle: \[P = 2(l + w)\] \[P = 2(15 + 3)\] \[P = 36\] You are correct.